ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 33 Driver, Kathy, University of the Witwatersrand, South Africa Zeros of Hypergeometric Functions The zeros of the Gauss hypergeometric function 2F1(-n, n+1;2n+2; z) are shown to cluster asymptotically as n tends to infinity on a simple circular arc. The application of this result to the asymptotic zero distribution of the associated Jacobi polynomials is also discussed.